OpenStudy (anonymous):

The opposite of the quotient of 24 and some number, x is equal to 12 less than the number, x. Choose the equation that could be used to find this number, x

OpenStudy (anonymous):

I'm not sure what is meant by "the opposite of the quotient". It could be the negative form, or the inverse of the fraction, or possibly something else entirely. The quotient is:\[\frac{24}{x}\]The opposite could be\[\frac{x}{24} \text{ or } -\frac{24}{x} \text{ or something else}\], and 12 less than the number x is \[x-12\]